# ITDR: An Instruction Tuning Dataset for Enhancing Large Language Models in Recommendations

### v1

@@ Abstract.
Large language models (LLMs) have demonstrated outstanding performance in natural language processing tasks. However, in the field of recommendation systems, due to the structural differences between user behavior data and natural language, LLMs struggle to effectively model the associations between user preferences and items. Although prompt-based methods can generate recommendation results, their inadequate understanding of recommendation tasks leads to constrained performance. To address this gap, in this work, we construct a sufficient instruction tuning dataset, ITDR, which encompasses 7 subtasks across two core root tasks—user-item interaction and user-item understanding. The dataset integrates data from 13 public recommendation datasets and is built using manually crafted standardized templates, comprising approximately 200,000 instances. Experimental results demonstrate that ITDR significantly enhances the performance of mainstream open-source LLMs such as GLM-4, Qwen2.5, Qwen2.5-Instruct and LLaMA-3.2 on recommendation tasks. Furthermore, we analyze the correlations between tasks and explore the impact of task descriptions and data scale on instruction tuning effectiveness. Finally, we perform comparative experiments against closed-source LLMs with substantial parameters. Our tuning dataset ITDR and the fine-tuned large recommendation models can be accessed at https://github.com/hellolzk/ITDR.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2026/08ccs: Information systems Recommender systems
@@ 1. Introduction
Recommendation systems, as a core technology to address the issue of information overload, have been driving the innovation of personalized user experiences in scenarios such as e-commerce, social media, and content platforms (REVIEWOFRECOMMENDERSYSTEMS). With the groundbreaking advancements in LLMs (zhao2025surveylargelanguagemodels), how to effectively leverage LLMs to enhance the performance of recommendation systems has become a cutting-edge topic of shared interest in both academia and industry (lin2025can). Among these efforts, instruction tuning techniques are regarded as a critical method to achieve this goal (RecLM; tallrec; Recommendationasinstructionfollowing; Parameter-efficientconversationalrecommender; lin2024rella; Data-efficientFine-tuningforLLM-basedRecommendation), owing to their ability to significantly improve the model’s understanding of specific user intents and task requirements.
However, current recommendation methods that rely on instruction-tuning of LLMs face a serious data bottleneck problem (wu2024surveylargelanguagemodels). First, traditional recommendation datasets primarily consist of ID-based explicit or implicit user-item interaction records (e.g., ratings or clicks), generally lacking structured task descriptions and diverse natural language instructions—precisely the prerequisites for effective fine-tuning of LLMs on recommendation tasks. Second, the diverse nature of recommendation tasks demands that training data comprehensively cover different types of task scenarios and user behavior patterns. The limitations of existing datasets severely constrain the generalization and adaptability of LLMs in recommendation scenarios, highlighting the urgency of constructing high-quality, diverse instruction datasets.
To address these challenges, we propose ITDR—a large-scale Instruction Tuning Dataset specifically tailored for Recommendation systems. ITDR integrates multi-dimensional data from 13 classic recommendation benchmarks (including MovieLens (MovieLens), Amazon Reviews (AmazonReviews2023), etc.), categorized into two core task classes—user-item interaction (UII) and user-item understanding (UIU)—and further subdivided into seven distinct subtasks. Each subtask is equipped with a meticulously designed task description to guide the LLM’s decision-making behavior. These subtasks comprehensively cover key recommendation scenarios such as rating prediction, sequential recommendation, and interest recognition, ensuring the dataset’s broad applicability. ITDR ultimately comprises nearly 200,000 high-quality instructions, and we commit to publicly releasing all datasets and fine-tuned models to foster continued progress in this field.
Based on the ITDR, we conducted systematic instruction tuning experiments on multiple open-source LLMs. Experimental results demonstrate that ITDR significantly enhances the performance of language models of varying scales across diverse recommendation tasks.
@@
1. The construction and open-sourcing of the ITDR benchmark dataset—a large-scale, task-comprehensive instruction fine-tuning dataset for recommendation systems, containing nearly 200,000 high-quality instructions across two root tasks and seven subtasks;
2. Rigorous experimental validation of ITDR’s effectiveness in enhancing the recommendation performance of LLMs, with thorough ablation studies providing critical references for future optimization and dataset expansion in recommendation-oriented LLMs.
@@ 2.1. Large Language Models for Recommendation
LLMs have demonstrated remarkable capabilities in natural language understanding and generation, achieving successful applications in various tasks such as text summarization and question answering. Leveraging their strengths in capturing complex linguistic patterns, researchers have begun exploring the potential of LLMs in recommender systems. Current research can be broadly categorized into two main approaches.The first approach directly employs LLMs as recommender systems by designing sophisticated prompt templates to harness their zero/few-shot (few-shot) learning capabilities. Representative works include: Chat-REC (Chat-rec), which transforms user profiles and interaction histories into prompts to enable conversational recommender systems; LLM4CDR (LLM4CRD) constructs context-aware prompts based on user’s purchase histories in source domains and cross-domain shared features to enhance cross-domain recommendation performance; and RecPrompt (Recprompt) develops an adaptive prompt framework for news recommendation that automatically optimizes prompt design.The second approach integrates LLMs with traditional recommender systems to augment their capabilities. Notable examples include Llmrec (Llmrec), which enhances graph representations of user-item interactions by incorporating LLM-generated features to alleviate data sparsity in implicit feedback scenarios, and LLM-CF (LLM-CF) distills the world knowledge and reasoning capabilities of LLMs into collaborative filtering frameworks, significantly improving recommendation effectiveness. While these methods can enhance LLMs‘ capabilities in recommendation tasks to some extent, they essentially rely on prompt engineering to elicit emergent abilities and align the models with recommendation objectives. However, due to the complexity of user behavior patterns and item features, as well as the scarcity of recommendation-related concepts in natural language text, larger models without fine-tuning do not perform well in some recommendation tasks (e.g., sequence recommendation) (Ischatgptagoodrecommender).
@@
2.2. Instruction Tuning for LLMs
@@
In this work, we focuse on optimizing the performance of LLMs in recommendation system-related tasks through instruction tuning techniques. We construct a meticulously designed dataset that fully incorporates the unique characteristics and practical application scenarios of recommendation tasks. Experimental results show that the adopted instruction tuning method can significantly enhance the model’s comprehensive performance across various recommendation tasks, validating the effectiveness and applicability of this technology in the field of recommendation systems.
@@ 3. ITDR: Instruction Tuning Dataset for Recommendation
In this section, we provide a detailed introduction to the ITDR dataset designed for instruction tuning tasks in recommender systems. Given the substantial resources required to construct a comprehensive instruction tuning dataset that covers diverse recommendation scenarios, we adopt the methodology of prior work (INTERS) by transforming existing recommendation datasets into an instructional format for efficient construction. Based on the characteristics of recommendation tasks, we categorize all tasks and their corresponding datasets into two root tasks: User-Item Interaction (UII) and User-Item Understanding (UIU). The complete taxonomy of tasks and their associated datasets is illustrated in Figure 1. Below, we elaborate on the specific definitions of these two root tasks, and the data construction process.
3.1. Root Task Definition
@@ 3.1.1. User-Item Interaction(UII)
In recommendation systems, user-item interactions refer to behavioral records between users and the content recommended by the system, typically including explicit or implicit signals such as clicks, views, purchases, and ratings. The core objective is to characterize the relationship between user preferences and item features, thereby providing a basis for personalized recommendations. In the recommendation process, modeling user-item interactions is a critical factor determining the accuracy of the recommendation system and the user experience. To this end, we employed four classical interaction recommendation tasks: Rating Prediction(RP), Top-K Recommendation(TKR), Cross-Domain Recommendation(CDR) and Next Item Recommendation(NIR), requiring the model to mine latent patterns in the interaction data and predict user’s interest levels in unexposed items.
@@ 3.1.2. User-Item Understanding(UIU)
In recommendation systems, user-item understanding refers to the in-depth feature mining from the item dimension. This process involves analyzing the inherent attributes of items (such as multi-dimensional features of movies, including genre, director, cast, plot synopsis, etc.) to construct accurate item profiles, thereby achieving personalized recommendation matching for target users. We design three tasks: User Attribute Prediction(UAP), Interest Recognition(IR) and Target User Identification(TUI), requiring the model to deeply comprehend item features and demonstrate better adaptability in personalized recommendation scenarios.
@@ 3.2. ITDR Construction
3. Example generation We use the task description as the ”instruction” part of the example and the template as the ”input” part of the example. For the UII task, the ground truth corresponding to each instruction can be directly obtained from the original dataset. As for the ”target user identification” and ”interest recognition” subtasks in the UIU task, since the original dataset does not include the corresponding true annotation data, we employ DeepSeek-V3 (Deepseek-v3) to generate the reference ground truth for these tasks.
After completing the above steps, we construct an ITDR instruction set comprising 2 root tasks and 7 subtasks, with a total scale of 195,065 instructions. Appendix A provides a detailed description of the root tasks and subtasks. Appendix B compiles comprehensive statistics and complete experimental results for ITDR. Appendix C offers specific illustrative examples for each subtask.
@@ 4.1. Backbone Models
We select four publicly available LLMs spanning different scales: GLM-4-9B (glm4), Qwen2.5-7B (qwen2.5) and its instruction-tuned variant Qwen2.5-7B-Instruct (qwen2.5), and LLaMA-3.2-3B (llama3.2). These models represent diverse architectures that have demonstrated strong performance on standard benchmarks and are widely applied in a variety of research studies.
@@ 4.2.2. Evaluation Metrics
Furthermore, building upon the methodology established in (INTERS), we define two composite metrics to holistically evaluate model performance on both user-item interaction and user-item understanding tasks. The construction methodology is as follows:
We first define the set of subtasks (ST) and the set of evaluation metrics for sequential recommendation tasks (M) as follows:
$ST=\{RT,TKR,CDR,NIR,UAP,IR,TUI\},$

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated outstanding performance in natural language processing tasks. However, in the field of recommender systems, due to the inherent structural discrepancy between user behavior data and natural language, LLMs struggle to effectively model the associations between user preferences and items. Although prompt-based methods can generate recommendation results, their inadequate understanding of recommendation tasks leads to constrained performance. To address this gap, we construct a comprehensive instruction tuning dataset, ITDR, which encompasses seven subtasks across two root tasks: user-item interaction and user-item understanding. The dataset integrates data from 13 public recommendation datasets and is built using manually crafted standardized templates, comprising approximately 200,000 instances. Experimental results demonstrate that ITDR significantly enhances the performance of mainstream open-source LLMs such as GLM-4, Qwen2.5, Qwen2.5-Instruct and LLaMA-3.2 on recommendation tasks. Furthermore, we analyze the correlations between tasks and explore the impact of task descriptions and data scale on instruction tuning effectiveness. Finally, we perform comparative experiments against closed-source LLMs with massive parameters. Our tuning dataset ITDR, the fine-tuned large recommendation models, all LoRA modules, and the complete experimental results are available at https://github.com/hellolzk/ITDR.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3785694isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems, as a core technology to mitigate information overload, have been catalyzing innovation in personalized user experiences across various domains, such as e-commerce, social media, and content platforms (REVIEWOFRECOMMENDERSYSTEMS). With the groundbreaking advancements in LLMs (zhao2025surveylargelanguagemodels), effectively leveraging LLMs to enhance recommendation performance has emerged as a pivotal research focus in both academia and industry (lin2025can). Among these advancements, instruction tuning techniques are regarded as a critical paradigm to achieve this goal (RecLM; tallrec; Recommendationasinstructionfollowing; Parameter-efficientconversationalrecommender; lin2024rella; Data-efficientFine-tuningforLLM-basedRecommendation), by virtue of their ability to significantly improve model understanding of specific user intents and task requirements.
However, current recommendation methods that leverage instruction tuning of LLMs face a severe data bottleneck (wu2024surveylargelanguagemodels). First, traditional recommendation datasets primarily consist of ID-based explicit or implicit user-item interaction records (e.g., ratings or clicks) and generally lack structured task descriptions and diverse natural language instructions–both of which are essential prerequisites for effective LLM fine-tuning in recommendation scenarios. Second, the heterogeneous nature of recommendation tasks necessitates that training data comprehensively cover multifaceted task scenarios and user behavior patterns. The limitations of existing datasets hinder the generalization and adaptability of LLMs in recommendation scenarios, highlighting the urgency of constructing high-quality, diverse instruction datasets.
To address these challenges, we propose ITDR–a large-scale Instruction Tuning Dataset specifically tailored for Recommendation systems. ITDR integrates multi-dimensional data from 13 classic recommendation benchmarks (including MovieLens (MovieLens), Amazon Reviews (AmazonReviews2023), etc.), categorized into two core task classes–user-item interaction (UII) and user-item understanding (UIU)–and further subdivided into seven distinct subtasks. Each subtask is equipped with a meticulously designed task description to guide the LLM’s decision-making process. These subtasks comprehensively cover key recommendation scenarios such as rating prediction, sequential recommendation, and interest recognition, ensuring the dataset’s broad applicability. ITDR ultimately comprises nearly 200,000 high-quality instructions, and we commit to publicly releasing the dataset and all fine-tuned models to foster continued progress in this field.
Based on ITDR, we conducted systematic instruction tuning experiments on multiple open-source LLMs. Experimental results demonstrate that ITDR significantly enhances the performance of language models of varying scales across diverse recommendation tasks.
@@
1. The construction of the ITDR benchmark dataset–a large-scale instruction tuning dataset for recommender systems with comprehensive task coverage, containing nearly 200,000 high-quality instructions across two root tasks and seven subtasks.
2. Rigorous experimental validation of ITDR’s effectiveness in enhancing the recommendation performance of LLMs, with thorough ablation studies providing critical insights for future optimization and dataset expansion in LLM-based recommender systems.
@@ 2.1. Large Language Models for Recommendation
LLMs have demonstrated remarkable capabilities in natural language understanding and generation, achieving successful applications in various tasks such as text summarization and question answering. Leveraging their strengths in capturing complex linguistic patterns, researchers have begun exploring the potential of LLMs in recommender systems. Current research can be broadly categorized into two main approaches.The first approach directly employs LLMs as recommender systems by designing sophisticated prompt templates to harness their zero/few-shot (gpt3) learning capabilities. Representative works include: Chat-REC (Chat-rec), which transforms user profiles and interaction histories into prompts to enable conversational recommender systems; LLM4CDR (LLM4CRD) constructs context-aware prompts based on user’s purchase histories in source domains and cross-domain shared features to enhance cross-domain recommendation performance; and RecPrompt (Recprompt) develops an adaptive prompt framework for news recommendation that automatically optimizes prompt design.The second approach integrates LLMs with traditional recommender systems to augment their capabilities. Notable examples include Llmrec (Llmrec), which enhances graph representations of user-item interactions by incorporating LLM-generated features to alleviate data sparsity in implicit feedback scenarios, and LLM-CF (LLM-CF) distills the world knowledge and reasoning capabilities of LLMs into collaborative filtering frameworks, significantly improving recommendation effectiveness. While these methods can enhance LLMs‘ capabilities in recommendation tasks to some extent, they essentially rely on prompt engineering to elicit emergent abilities and align the models with recommendation objectives. However, due to the complexity of user behavior patterns and item features, as well as the scarcity of recommendation-related concepts in natural language text, larger models without fine-tuning do not perform well in some recommendation tasks (e.g., sequence recommendation) (Ischatgptagoodrecommender).
@@
2.2. Fine-tuning for LLMs
2.2.1. Instruction tuning
@@
In terms of technical implementation, instruction tuning necessitates the reformulation of training data into input-output pairs $(x,y)$, where $x$ denotes human instructions expressed in natural language and $y$ represents the corresponding model responses. This format not only explicitly encodes the semantic descriptions of target tasks but also requires the systematic transformation of raw training samples into natural language forms aligned with human expression, thereby constructing semantically complete and logically coherent instructional contexts. Building upon this foundation, Large Language Models (LLMs) can be fine-tuned using these $(x,y)$ pairs based on the autoregressive language modeling objective:
(1) $\max_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{t=1}^{|y|}\log P_{\theta}\left(y_{t}|x,y_{<t}\right)\right].$
where $\theta$ represents the set of trainable parameters of the large language model (LLM), $y_{t}$ denotes the $t$-th token in the target sequence $y$, $y_{<t}$ refers to the sequence of tokens generated prior to time step $t$, and $\mathcal{D}$ denotes the empirical distribution over the training dataset. This objective maximizes the expected log-likelihood of generating each token autoregressively, conditioned on the input and previous outputs.
2.2.2. Fine-tuning with Low-rank Adaptation (LoRA)
Full fine-tuning of large language models typically demands substantial computational resources. To address this challenge, Low-Rank Adaptation (hu2021loralowrankadaptationlarge) (LoRA) has been proposed as an efficient parameter-efficient fine-tuning approach. The core idea of LoRA is to inject learnable low-rank matrices into specific layers of the Transformer architecture to approximate the weight updates required during fine-tuning of the pre-trained model. Specifically, for a pre-trained weight matrix ${W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, its update $\Delta{W}$ is factorized as the product of two low-rank matrices:
$\Delta{W}={B}{A},$
where ${B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and ${A}\in\mathbb{R}^{r\times d_{\text{in}}}$ are trainable upward and downward projection matrices, respectively, and the rank $r$ is significantly smaller than both $d_{\text{in}}$ and $d_{\text{out}}$. This low-rank parameterization drastically reduces the number of trainable parameters and enhances adaptation efficiency.
In practice, LoRA is primarily applied to the query and value projection matrices within the multi-head attention mechanism of Transformer layers. Given an input vector h{h} to such a linear transformation, the modified output ${h}^{\prime}$ under LoRA is expressed as:
(2) ${h}^{\prime}=\left({W}+\frac{\alpha}{r}\Delta{W}\right){h}={W}{h}+\frac{\alpha}{r}{B}{A}{h},$
where the original weights W{W} remain frozen during fine-tuning, and $\alpha$ is a tunable scaling factor that modulates the influence of the low-rank update relative to the base weights.
This strategy enables effective task-specific adaptation with minimal architectural modification, circumventing the need for full retraining of all model parameters while preserving strong downstream performance–all at a fraction of the computational cost.
In this work, we enhance LLM performance on recommendation tasks via instruction tuning with a carefully designed dataset that captures key characteristics and real-world scenarios of recommender systems. Experiments show the dataset effectively supports tuning and significantly boosts performance across multiple tasks, highlighting its value and applicability.
@@ 3. ITDR: Instruction Tuning Dataset for Recommendation
In this section, we provide a detailed introduction to ITDR. Given the substantial resources required to construct a comprehensive instruction tuning dataset that covers diverse recommendation scenarios, we adopt the methodology of prior work (zhu2024intersunlockingpowerlarge) by transforming existing recommendation datasets into an instructional format for efficient construction. Based on the characteristics of recommendation tasks, we categorize all tasks and their corresponding datasets into two root tasks: User-Item Interaction (UII) and User-Item Understanding (UIU). The complete taxonomy of tasks and their associated datasets is illustrated in Figure 1. Below, we elaborate on the specific definitions of these two root tasks and their corresponding subtasks, along with the data construction process.
3.1. Task Definition
@@ 3.1.1. User-Item Interaction(UII)
In recommendation systems, user-item interactions refer to behavioral records between users and the content recommended by the system, typically including explicit or implicit signals such as clicks, views, purchases, and ratings. The core objective is to characterize the relationship between user preferences and item features, thereby providing a basis for personalized recommendations. In the recommendation process, modeling user-item interactions is a critical factor determining the accuracy of the recommendation system and the user experience. To this end, we employed four classical interaction recommendation tasks: Rating Prediction(RP), Top-K Recommendation(TKR), Cross-Domain Recommendation(CDR) and Next Item Recommendation(NIR), requiring the model to mine latent patterns in the interaction data and predict user’s interest levels in unexposed items. The following paragraphs provide detailed introductions to each subtask under the root task UII and list the source datasets we selected.
Rating Prediction (RP)
Top-K Recommendation (TKR)
Cross-Domain Recommendation (CDR)
Table 1. Selection strategies for source and target domains in CDR task.
Source Target
Beauty Fashion
CDs & Vinyl Video Games
Appliances Industrial & Scientific
Home & Kitchen Grocery & Gourmet Food
Next Item Recommendation (NIR)
@@ 3.1.2. User-Item Understanding(UIU)
In recommendation systems, user-item understanding refers to the in-depth feature mining from the item dimension. This process involves analyzing the inherent attributes of items (such as multi-dimensional features of movies, including genre, director, cast, plot synopsis, etc.) to construct accurate item profiles, thereby achieving personalized recommendation matching for target users. We design three tasks: User Attribute Prediction(UAP), Interest Recognition(IR) and Target User Identification(TUI), requiring the model to deeply comprehend item features and demonstrate better adaptability in personalized recommendation scenarios. The following paragraphs provide detailed introductions to each subtask under the root task UIU and list the source datasets we selected.
User Attribute Prediction (UAP)
Interest Recognition (IR)
The primary objective of interest recognition is to analyze user‘s historical interaction data to identify and summarize their preferences, thereby enabling more accurate personalized services. For this subtask, we select the following datasets as data sources: Anime Dataset 20231, Amazon Books Reviews2, and Steam (steam).
Target User Identification (TUI)
@@ 3.2. ITDR Construction
3. Example generation We use the task description as the ”instruction” part of the example and the template as the ”input” part of the example. For the UII task, the ground truth corresponding to each instruction can be directly obtained from the original dataset. As for the ”target user identification” and ”interest recognition” subtasks in the UIU task, since the original dataset does not include the corresponding true annotation data, we employ DeepSeek-V3 (Deepseek-v3) to generate the reference ground truth for these tasks. Following label generation, we conduct random sampling and manual verification: specifically, for each dataset under both subtasks, we randomly sample 100 instances for human review. Our evaluation confirms that the outputs generated by DeepSeek-V3 are sufficiently accurate and fluent to serve as reliable ground-truth labels.
After completing the above steps, we construct an ITDR instruction set comprising 2 root tasks and 7 subtasks, with a total scale of 195,065 instructions. Appendix A compiles comprehensive statistics and complete experimental results for ITDR. Appendix B offers specific illustrative examples for each subtask.
3.3. Notations and Terminologies
For clarity in describing ITDR and its associated evaluation metrics, we begin by standardizing the key notations used in this work. Table 2 provides a summary of the primary symbols and their meanings.
Table 2. Key notations.
Symbol Definition
UII User-Item Interaction
UIU User-Item Understanding
RP Rating Prediction
TKR Top-K Recommendation
CDR Cross-Domain Recommendation
NIR Next Item Recommendation
UAP User Attribute Prediction
IR Interest Recognition
TUI Target User Identification
STST Set of subtasks
$M$ Set of evaluation metrics for sequential recommendation tasks
$M^{\prime}$ Set of evaluation metrics for IR and TUI tasks
$D_{st}$ Set of all subtasks data
$D_{UAP}$ Set of UAP task
pUIIpUII Performance of user-item interaction
pUIUpUIU Performance of user-item understanding
@@ 4.1. Backbone Models
We select four open-source LLMs with parameter counts ranging from 3B to 9B.
$\bullet$ LLaMA-3.2-3B (llama3.2) is a lightweight large language model developed by Meta AI. It supports a context length of up to 128K tokens and is well-suited for natural language understanding and generation tasks in edge and mobile computing scenarios, including text summarization, multilingual dialogue, and tool invocation [^3] https://www.modelscope.cn/models/LLM-Research/Llama-3.2-3B.
$\bullet$ Qwen2.5-7B (qwen2.5) and its instruction-tuned variant [^4] https://www.modelscope.cn/models/Qwen/Qwen2.5-7B-Instruct are large language models developed by the Tongyi Qianwen team at Alibaba Cloud. These models are pretrained on a high-quality multilingual corpus exceeding 18 trillion tokens and support a context length of up to 128K tokens. They are highly capable of handling diverse tasks such as code generation, mathematical reasoning, and structured output generation, delivering robust performance in both general-purpose AI assistant and enterprise settings [^5] https://www.modelscope.cn/models/Qwen/Qwen2.5-7B.
$\bullet$ GLM-4-9B (glm4) is a large language model released by Zhipu AI. It is pretrained on 10 trillion tokens of high-quality multilingual corpus, and supports a context window of 8K tokens, making it suitable for general-purpose natural language understanding and generation tasks [^6] https://www.modelscope.cn/models/ZhipuAI/glm-4-9b.
The aforementioned models represent diverse architectures, have demonstrated strong performance on standard benchmarks and are widely applied in a variety of research studies.
@@ 4.2.2. Evaluation Metrics
Furthermore, building upon the methodology established in (zhu2024intersunlockingpowerlarge), we define two composite metrics to holistically evaluate model performance on both user-item interaction and user-item understanding tasks. The construction methodology is as follows: We first define the set of subtasks (ST) and the set of evaluation metrics for sequential recommendation tasks (M) as follows:
$ST=\{RP,TKR,CDR,NIR,UAP,IR,TUI\},$